the length of a rectangular width is 10 ft less than the length and the perimeter of 44 ft what should be the length
Answer by checkley77(12569)

W=L-10
2L+2W=P
2L+2(L-10)=44
2L+2L-20=44
4L=44+20
4L=64
L=64/4
L=16 ANS. FOR THE LENGTH.
W=16-10
W=6 ANS. FOR THE LENGTH.
PROOF:
2+16+2*6=44
32+12=44
44=44